OpenStudy (anonymous):

Does anyone know where I would start with this and what method I would use? Q. Suppose det A = 3, B = -2/3 and C = 1/2. Calculate det (A^2B^-1C).

OpenStudy (anonymous):

Am I supposed to use detA^-1 formula = 1/detA so then my answer becomes det (3^2 * 1 / -2/3 * 1/2 ) = -27/4

OpenStudy (phi):

yes for matrices A and B det(A*B)= det(A)* det(B) also, det(A_inv) = 1/det(A) so you get det(A)*det(A)*det(B_inv)*det© 3*3* * -3/2 * ½ = -27/4 which is what you got.

OpenStudy (anonymous):

thank you! for b) det (A^-1B^2C^-1)^-1 my workings are: det ( 1 / ( 1/3*-2^2/3* 1/1/2) = 17/9 Is this also correct?

OpenStudy (anonymous):

sorry answer should be 27/8

OpenStudy (phi):

A= 3 , B= -⅔, C= ½ <--oops, C is ½ not 2 ⅓ * (-⅔ * -⅔) * 2 = 8/27 invert the whole thing to get 27/8

OpenStudy (anonymous):

Thanks :)

OpenStudy (phi):

yw